The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                 http://www.astro.uni-wuerzburg.de/~grebel/MCNews.html

No. 9 Feb 11, 1997

Table of Contents

News and Views............................1 Refereed Paper Abstracts (12)......2 Invited Review Abstract (1)........10

News and Views

Proposal for an IAU Symposium on the Magellanic Clouds

As announced in the previous Magellanic Clouds Newsletters, a symposium on the Magellanic Clouds is being planned and will be proposed to the IAU in April. A scientific organization committee is being assembled by You-Hua Chu, Nick Suntzeff, and Jim Hesser. Progress on the preparation will be posted at the web page http://www.astro.uiuc.edu/mcnews/MC IAUsymp.html

Widening the Scope of the Magellanic Clouds Newsletter?

It has been suggested that the scope of the Magellanic Clouds Newsletter be widened to include other similar irregular galaxies. Through comparison between the Magellanic Clouds and other irregular galaxies, we may better understand the structure and evolution of the Magellanic Clouds. Please send comments and suggestions to mcnews@astro.uiuc.edu. Thanks.

Abstracts of Refereed Papers

A radio continuum study of the Magellanic Clouds. VI. Discrete sources common to radio and X-ray surveys of the Magellanic Clouds

M.D. Filipović^{1,2,3}, W. Pietsch^{3}, R.F. Haynes^{2,1}, G.L. White^{1}, P.A. Jones^{1}, R. Wielebinski^{4}, U. Klein^{5}, K. Dennerl^{3}, P. Kahabka^{3,6} and J.S. Lazendić^{7}

1 University of Western Sydney, Nepean, P.O. Box 10, Kingswood, NSW 2747 Australia 2 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121 Australia 3 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85740 Garching, Germany 4 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany 5 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 6 Astronomical Institute and Center for High Energy Astrophysics, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands ; 7 Katedra za Astronomiju, Univerzitet u Beogradu, Studentski trg 13, 11000 Beograd, Serbia

A comparison of Parkes radio surveys and the X-ray ROSAT All-Sky Survey (RASS) of the Large Magellanic Cloud (LMC) has found 71 discrete sources of both radio and X-ray emission. These 71 sources are mainly supernova remnants (SNRs) and SNR candidates (36) and background sources (27). For six of the sources we have no proposed identification, and two others are H ii regions. A source intensity comparison of the radio and X-ray sources shows very little correlation, but we note that the strongest SNRs at both radio and X-ray frequencies are young SNRs from Population I. Six new LMC SNR candidates are proposed. From the radio flux density of the SNRs we have estimated the SNR birth rate to be one every 100 (±20) years and the star formation rate (SFR) to be 0.7 (±0.2) M_{⊙} yr^{−1}.

A similar comparison was undertaken for the Small Magellanic Cloud (SMC), but instead of the RASS we used a roster of pointed observations made with the ROSAT Position Sensitive Proportional Counter (PSPC). This comparison resulted in 27 sources in common between Parkes radio and ROSAT PSPC surveys. Two new SMC sources are proposed for SNR candidates. The SMC SNR birth rate was estimated to be one every 350 (±70) years and the SFR was estimated to be 0.15 (±0.05) M_{⊙} yr^{−1}. Submitted to: Astronomy & Astrophysics

For preprints, contact mfilipov@atnf.csiro.au

The X-ray shadow in the south-east of the LMC

M.J. Blondiau^{1}, J. Kerp^{1}, U. Mebold^{1} and U. Klein^{1}

1 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany

We analyze mosaicked ROSAT PSPC pointings towards the Large Magellanic Cloud (LMC). Focusing on the area south-east of 30 Doradus roughly at the position of RA(2000) = 5^{h}40^{m} and DEC(2000) = −70^{∘}, we verify the existence of an X-ray absorption feature. The ROSAT data imply that X-ray-attenuating material is located in front of a diffuse X-ray source that extends from the eastern boundary of the LMC to west of the giant molecular cloud south of 30 Dor. The difference of the absorbing hydrogen column densities, derived from the ROSAT data, between the X-ray-bright “wedge” region and the dark “shadow” region is about N_{H} = (7.0 ± 1.8) · 10^{21} cm^{−2}. This column density difference is consistent with the value derived from the IRAS 100 μm emission map of this region, which leads to total hydrogen column densities of N_{Wedge} = (5.2 ± 0.9) · 10^{21} cm^{−2} towards the X-ray-bright region, and N_{Shadow} = (11.1 ± 2.2) · 10^{21} cm^{−2} towards the X-ray-dark regions. The agreement between these values derived from the ROSAT and IRAS data suggests that part of the extended diffuse X-ray emission is located on the far side of the giant molecular cloud.

This kind of 3-D structure is independently verified by radio continuum observations. The X-ray attenuating gas is positionally associated with the polarized radio continuum emission of a giant magnetic loop emerging from the plane of the LMC. Faraday depolarization implies that this magnetized loop penetrates deep into – or beyond – the LMC, possibly commencing in the X-ray-emitting plasma of the LMC. Accepted by: Astronomy & Astrophysics

For preprints, contact blondiau@astro.uni-bonn.de
Also available from the URL http://www.astro.uni-bonn.de/~blondiau

Supernova Remnants in OB Associations

Y.-H. Chu^{1}

1 Astronomy Department, University of Illinois, 1002 W Green St., Urbana, IL 61801, USA

We use the sample of supernova remnants (SNRs) in OB associations in the Large Magellanic Cloud (LMC) to study how interstellar environments affect the physical properties of SNRs. Some of these SNRs have been identified by the three classical SNR signatures - bright X-ray emission, nonthermal radio emission, and enhanced [S ii]/Hα ratio; the others have been diagnosed by bright Xray emission from superbubbles. We have obtained long-slit echelle observations of these 14 SNRs and their surroundings to search for high-velocity shocked gas in the SNRs and to examine the structure of the ambient interstellar medium. Not all SNRs show detectable high-velocity (ΔV ≥ 100 km s^{−1}) gas. SNRs in H ii regions all show classical SNR signatures and high-velocity gas, while SNRs in superbubbles often show only diffuse X-ray emission but not the other SNR signatures. Using ROSAT X-ray observations of two superbubbles, N51D and N44, we illustrate quantitatively that SNR shocks in the interiors of superbubbles can produce bright, detectable X-ray emission for ROSAT and Einstein detectors, but not enough Hα emission for echelle observations using a 4-m class telescope. The wide variety of SNR properties can be explained by the differences in their interstellar environments. Accepted by: the Astronomical Journal (May 1997 issue) For preprints, contact chu@astro.uiuc.edu

Also available from the URL http://www.astro.uiuc.edu/~chu/preprints/lmcobsnr.ps

Uncovering A Supernova Remnant Hidden Near LMC X-1

Y.-H. Chu^{1}, R.C. Kennicutt^{2}, S.L. Snowden^{3}, R.C. Smith^{4}, R.M. Williams^{1}, D.J. Bomans^{1}

1 Astronomy Department, University of Illinois, Urbana, IL 61801 2 Astronomy Department, University of Arizona, Tucson, AZ 85721 3 Goddard Space Flight Center, NASA, Greenbelt, MD 20771 4 Astronomy Department, University of Michigan, Ann Arbor, MI 48109-1090

In a kinematic study of LMC H ii regions, we serendipitously found a fast-expanding (velocity offsets up to ∼200 km s^{−1}) shell in the northern lobe of the luminous H ii region N159. The expansion pattern and the [S ii]/Hα ratio of the shell are consistent with those seen in supernova remnants (SNRs). Confirmation of this SNR candidate in radio and X-ray wavelengths has been hampered by thermal emission of the luminous H ii region N159 itself and scattered X-rays from the nearby luminous source LMC X-1. Recent ROSAT observations finally resolved N159 from LMC X-1, thus confirming the identification of this as a newly identified SNR, which we designate as SNR 0540−697. Accepted by: Publications of the Astronomical Society of the Pacific, May 1997 issue For preprints, contact chu@astro.uiuc.edu

or from the URL http://www.astro.uiuc.edu/~chu/preprints/N159.ps

Spectroscopic study of the planetary nebulae with Wolf-Rayet nuclei in the Magellanic Clouds

M. Peña^{1}, M. T. Ruiz^{2}, and S. Torres-Peimbert^{1}

1 Instituto de Astronomía, Universidad Nacional Autónoma de México, Apdo. Postal 70 264, México D.F. 045 10, México 2 Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

UV and optical spectrophotometric data of planetary nebulae with WR nuclei in the Magellanic Clouds are presented. Analysis of the nebular characteristics shows that the ionized gas, in all the objects, is extremely C-rich, probably due to contamination with freshly-made C from stellar nucleosynthesis. Very high electron densities are found (n_{e} ≥ 10^{4} cm^{−3}), which are interpreted as an indication of youth. The central stars are C-rich, with spectral types in the range from WC 4/5 to WC 8. The spectral type distribution is totally different from that found for galactic planetary nebulae with WR nuclei. Estimates of the stellar visual magnitudes and temperatures are presented. Stellar temperatures over 60 000 K were derived for all the objects, with some cases exceeding 100 000 K. These high temperatures, however, do not produce high excitation nebulae. Accepted by: Astronomy & Astrophysics

For preprints, contact miriam@astroscu.unam.mx

Ultraviolet Imaging Telescope (UIT) Observations of the Small Magellanic Cloud

Robert H. Cornett ^{1,2}, Michael R. Greason ^{1,2}, Jesse K.Hill ^{1,2}, Joel Wm. Parker^{3}, William H. Waller^{1,2}, Ralph C.Bohlin^{4}, Kwang-Peng Cheng^{5}, Susan G. Neff^{2}, Robert W. O’Connell^{6}, Morton S. Roberts^{7}, Andrew M. Smith ^{2}, and Theodore P. Stecher ^{2}

1 Hughes STX Corporation 2 Laboratory for Astronomy and Solar Physics, Code 681, GSFC, Greenbelt MD 20771 3 Southwest Research Institute, 1050 Walnut St., Suite 426, Boulder CO 80302 4 STScI, Homewood Campus, Baltimore MD 21218 5 Cal. State Univ., Fullerton, Dept. of Physics, 800 N. State College Ave., Fullerton, CA 92634 6 Univ. of Virginia, Astronomy Dept., P.O. Box 3818, Charlottesville VA 22903-0818 7 NRAO, 520 Edgemont Rd., Charlottesville, VA 22903-2475

A mosaic of four UIT far-UV (FUV) (λ_{eff} = 1620Å) images, with derived stellar and HII region photometry, is presented for most of the Bar of the SMC. The UV morphology of the SMC’s Bar shows that recent star formation there has left striking features including: a) four concentrations of UV-bright stars spread from northeast to southwest at nearly equal (∼30 arcmin=0.5 kpc) spacings; b) one of the concentrations, near DEM 55, comprises a well-defined 8-arcmin diameter ring surrounded by a larger Hα ring, suggestive of sequential star formation.

FUV PSF photometry is obtained for 11,306 stars in the FUV images, resulting in magnitudes m(1620). We present a FUV luminosity function for the SMC bar, complete to m(1620)∼14.5. Detected objects are well correlated with other SMC Population I material; of 711 Hα emission-line stars and smallnebulae within the UIT fields of view, 520 are identified with FUV sources. The FUV photometry is compared with available ground-based catalogs of supergiants, yielding 191 detections of 195 supergiants with spectral type earlier than F0 in the UIT fields. The (m(1620)−V) color for supergiants is a sensitive measure of spectral type. The bluest observed colors for each type agree well with colors computed from unreddened Galactic spectral atlas stars for types earlier than about A0; for later spectral types the observed SMC stars range significantly bluer, as predicted by comparison of low-metallicity and Galactic-composition models. Redder colors for some stars of all spectral types are attributed to the strong FUV extinction arising from even small amounts of SMC dust. Internal SMC reddenings are determined for all catalog stars. All stars with E(B−V)>0.15 are within regions of visible Hα emission.

FUV photometry for 42 Hα-selected HII regions in the SMC Bar is obtained for stars and for total emission (as measured in HII-region-sized apertures). The flux-weighted average ratio of total to stellar FUV flux is 2.15; consideration of the stellar FUV luminosity function indicates that most of the excess total flux is due to scattered FUV radiation, rather than stars fainter than m(1620)=14.5. Both stellar and total emission are well by Kennicutt and Hodge (1986; hereafter KH), yielding FUV/Hα flux ratios that are consistent with models of SMC metallicity, ages from 1-5 Myr, and moderate (E(B−V)=0.0–0.1 mag) internal SMC extinction. Accepted by: The Astronomical Journal (March 1997 issue)

For preprints, contact cornett@uit.gsfc.nasa.gov

Also available from the URL http://xxx.lanl.gov

Ultraviolet Photometry of Stars in the Compact Cluster R136 in the Large Magellanic Cloud

Deidre A. Hunter^{1}, William D. Vacca^{2}, Philip Massey^{3}, Roger Lynds^{3}, and Earl J. O’Neil^{3}

1 Lowell Observatory, 1400 West Mars Hill Road, Flagstaff, Arizona 86001, USA 2 Institute for Astronomy, 2680 Woodlawn Drive, Honolulu, Hawaii 96822, USA 3 National Optical Astronomy Observatories, PO Box 26732, Tucson, Arizona 85726, USA

We present ultraviolet photometry of stars in the Large Magellanic Cloud star cluster R136. The data were obtained with the refurbished Wide Field/Planetary Camera on the Hubble Space Telescope through an ultraviolet filter centered at 1750 Å (F170W). These data, combined with optical HST observations presented previously, are used to explore the hot, luminous stellar population within the cluster. The first question we wanted to address was whether the ultraviolet could place better constraints on differential reddening and the degree of coevality within the cluster than optical photometry alone. The stellar sequence in the color magnitude diagram using the F170W−F555W color is broader than it was in the optical and is broader than would be expected from photometric uncertainties. Although there could be a modest age spread among the massive stars, a large part of the intrinsic spread in the color-magnitude diagram is likely to be due to reddening differences. There is also some evidence that stars in the outer parts of the cluster are affected by a larger range in reddening than those in the smaller region of the cluster core. The second question we addressed is whether ultraviolet photometry in combination with optical photometry can distinguish the most massive stars more readily than is possible with the optical colors. The F170W−F555W color is better at separating B supergiants, of which there is only one in R136, from comparably bright O and Wolf-Rayet stars. However, spectral classifications of the stars within R136 will be necessary to properly identify stellar types of the rest of the stars and to disentangle age and reddening effects. Accepted by: The Astronomical Journal For preprints, contact dah@lowell.edu

Spectral Classification of the 30 Doradus Stellar Populations

Nolan R. Walborn^{1,2,3} and J. Chris Blades^{1,2}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 Visiting Observer, Anglo-Australian Telescope 3 Visiting Astronomer, Cerro Tololo Inter-American Obs., National Optical Astronomy Obs.

An optical spectral classification study of 106 OB stars within the 30 Doradus Nebula has sharpened the description of the spatial and temporal structures among the associated clusters. Five distinct stellar groups are recognized: (1) the central early-O (Carina phase) concentration which includes R136; (2) a younger (Orion phase) population to the north and west of R136 containing heavily embedded early-O dwarfs and IR sources, the formation of which was likely triggered by the central concentration; (3) an older population of late-O and early-B supergiants (Scorpius OB1 phase) throughout the central field, whose structural relationship if any to the younger groups is unclear;

(4) a previously known, older still compact cluster 3′ northwest of R136 containing A- and M-type supergiants (h and χ Persei phase), which evidently affects the nebular dynamics substantially; and (5) a newly recognized Sco OB1-phase association surrounding the recently discovered Luminous Blue Variable R143 in the southern part of the Nebula. The intricacy of this region and the implications for the interpretation of more distant starbursts are emphasized. The evidence is that the formation of the 30 Dor stellar content was neither instantaneous nor continuous, but rather that it corresponds to discrete events at different epochs.

The average difference between the derived and calibration absolute visual magnitudes of the stars is 0.05, indicating that the classification, calibration, and adopted distance modulus (V_{0} − M_{V} = 18.6) are accurate. For 70 of the stars, either that difference is ≤ 0.6 mag, or they are subluminous dwarfs or superluminous supergiants. A number of astrophysically interesting objects have been isolated for further investigation. Surprisingly in view of the presence of several O3 supergiants, the mid-Of star R139 is identified as the most massive object in this sample; it is located well along the 120 M_{⊙} track, very near the Humphreys-Davidson limit, and it is likely an immediate LBV precursor. This work can and should be extended in three ways: (1) higher-resolution and –S/N observations of many of the stars with larger ground-based telescopes for quantitative analysis; (2) ground-based spectral classification of the numerous additional accessible stars in the field; and (3) spatially resolved spectral classification of compact multiple systems with the Hubble Space Telescope. Submitted to: The Astrophysical Journal Supplements For preprints, contact walborn.stsci.edu

The initial conditions of young globular clusters in the LMC

Simon P. Goodwin

Astronomy Centre, University of Sussex, Falmer, Brighton, BN1 9QH, UK

N-body simulations are used to model the early evolution of globular clusters. These simulations include residual gas not turned into stars which is expelled from the globular cluster by the actions of massive (> 8 M_{⊙}) stars. The results of these simulations are compared to observations of 8 LMC globular clusters less than 100 Myr old. These observations are used to constrain the initial conditions that may have produced these clusters. It is found that the observations can be accounted for in a model where the globular clusters form from proto-cluster clouds similar to Plummer models with length scales in the range 1 < R_{S}/pc< 3 where the star formation efficiency varies between 25% and 60%. Using these derived initial conditions the survivability of these clusters in both the Galaxy and the LMC is assessed. If the slope of the initial mass function is around α = 2.35 then 2 or 3 of these clusters may be able to survive for a Hubble time even in the Galactic halo. In this case these clusters may represent young versions of the Galactic globular cluster population which was severely depleted by the destruction of many of its original members. In the case where α = 1.50, however, none of these clusters would be expected to survive for more than a few Gyr at most, even within the LMC. Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact spg@star.maps.susx.ac.uk
Also available by anonymous ftp at star.maps.susx.ac.uk/pub/papers/spg/lmc.ps.Z
or from the URL http://star-www.maps.susx.ac.uk/people/spg.html

The ellipticities of Galactic and LMC globular clusters

Simon P. Goodwin

Astronomy Centre, University of Sussex, Falmer, Brighton, BN1 9QH, UK

The correlations between the ellipticity and the age and mass of LMC globular clusters are examined and both are found to be weak. It is concluded that neither of these properties are mainly responsible for the observed differences in the LMC and Galactic globular cluster ellipticity distributions. Most importantly, age cannot be the primary factor in the LMC-Galaxy ellipticity differences, even is there is a relationship, as even the oldest LMC clusters are more elliptical than their Galactic counter-parts.

The strength of the tidal field of the parent galaxy is proposed as the dominant factor in determining the ellipticities of that galaxy’s globular clusters. A strong tidal field rapidly destroys velocity anisotropies in initially triaxial, rapidly rotating elliptical globular clusters. A weak tidal field, however, is unable to remove these anisotropies and the clusters remain close to their initial shapes. Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact spg@star.maps.susx.ac.uk
Also available by anonymous ftp at star.maps.susx.ac.uk/pub/papers/spg/ellip.ps.Z
or from the URL http://star-www.maps.susx.ac.uk/people/spg.html

Metallicity Effects on the Cepheid Extragalactic Distance Scale from EROS photometry in LMC and SMC

D.D. Sasselov^{1,2}, J.P. Beaulieu^{1}, C. Renault^{3}, P. Grison^{1}, R. Ferlet^{1}, A. Vidal-Madjar^{1},

E. Maurice^{4}, L. Prévot^{4}, E. Aubourg^{3}, P. Bareyre^{3}, S. Brehin^{3}, C. Coutures^{3}, N. Delabrouille^{3}, J. de Kat^{3}, M. Gros^{3}, B. Laurent^{3}, M. Lachièze-Rey^{3}, E. Lesquoy^{3}, C.

Magneville^{3}, A. Milsztajn^{3}, L. Moscoso^{3}, F. Queinnec^{3}, J. Rich^{3}, M. Spiro^{3}, L. Vigroux^{3}, S. Zylberajch^{3}, R. Ansari^{5}, F. Cavalier^{5}, M. Moniez^{5}, C. Gry^{6}, J. Guibert^{7},

O. Moreau^{7}, and F. Tajhmady^{7} (The EROS collaboration).

1 Institut d’Astrophysique de Paris, CNRS, 98bis Boulevard. Arago, 75014 Paris, France 2 Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA 3 CEA, DSM/DAPNIA, Centre d’études de Saclay, 91191 Gif-sur-Yvette, France 4 Observatoire de Marseille, 2 place Le Verrier, 13238 Marseille 03, France 5 Laboratoire de l’Accélérateur Linéaire IN2P3, Centre d’Orsay, 91405 Orsay, France 6 Laboratoire d’Astronomie Spatiale CNRS, Traversée du siphon, les trois lucs, 13120 Marseille, France 7 Centre d’Analyse des Images de l’Institut National des Sciences de l’Univers, CNRS Observatoire de Paris, 61 Avenue de l’Observatoire, 75014 Paris, France

This is an investigation of the period-luminosity relation of classical Cepheids in samples of different metallicity. It is based on 481 Cepheids in the Large and Small Magellanic Clouds from the CCD BRobservations data set of the French EROS microlensing project. The data set is complete and provides an excellent basis for a differential analysis between LMC and SMC. In comparison to previous studies of effects on the PL-relation, the EROS data set offers extremely well-sampled light curves and wellfilled instability strips. This allows reliable separation of Cepheids pulsating in the fundamental and the first overtone mode and derivation of differential reddening.

Our main result concerns the determination of distances to galaxies which are inferred by using the LMC as a base and using two-color photometry to establish the amount of reddening. We find a zero-point offset between SMC and LMC which amounts to a difference between inferred and true distance modulus of 0.14±0.06 mag in the V I_{c} system. The offset is exactly the same in both sets of PL-relations − of the fundamental and of the first overtone mode Cepheids. No effect is seen on the slopes of the PL-relations, although the fundamental and the first overtone mode Cepheids have different PL slopes. We attribute the color and the zero-point offset to the difference in metallicity between the SMC and LMC Cepheids. A metallicity effect of that small a magnitude still has important consequences for the inferred Cepheid distances and the determination of H_{0}. When applied to recent estimates based on HST Cepheid observations, our metallicity dependence makes the low-H_{0} values (Sandage et al. 1994) higher and the high-H_{0} values (Freedman et al. 1994b) lower, thus bringing those discrepant estimates into agreement near H_{0} ∼ 70 km s^{−1} Mpc^{−1}. Accepted by: Astronomy and Astrophysics

For preprints, contact beaulieu@astro.rug.nl, sasselov@cfassp36.harvard.edu
Also available from the URL http://xxx.lanl.gov/ps/astro-ph/9612216

The effect of metallicity on the Cepheid distance scale and its implications for the Hubble constant (H_{0}) determination

J.P. Beaulieu^{1}, D.D. Sasselov^{2,1}, C. Renault^{3}, P. Grison^{1}, R. Ferlet^{1}, A. Vidal-Madjar^{1},

E. Maurice^{4}, L. Prévot^{4}, E. Aubourg^{3}, P. Bareyre^{3}, S. Brehin^{3}, C. Coutures^{3}, N. Delabrouille^{3}, J. de Kat^{3}, M. Gros^{3}, B. Laurent^{3}, M. Lachièze-Rey^{3}, E. Lesquoy^{3}, C.

Magneville^{3}, A. Milsztajn^{3}, L. Moscoso^{3}, F. Queinnec^{3}, J. Rich^{3}, M. Spiro^{3}, L. Vigroux^{3}, S. Zylberajch^{3}, R. Ansari^{5}, F. Cavalier^{5}, M. Moniez^{5}, C. Gry^{6}, J. Guibert^{7},

O. Moreau^{7}, and F. Tajhmady^{7} (The EROS collaboration).

1 Institut d’Astrophysique de Paris, CNRS, 98bis Boulevard. Arago, 75014 Paris, France 2 Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA 3 CEA, DSM/DAPNIA, Centre d’études de Saclay, 91191 Gif-sur-Yvette, France 4 Observatoire de Marseille, 2 place Le Verrier, 13238 Marseille 03, France 5 Laboratoire de l’Accélérateur Linéaire IN2P3, Centre d’Orsay, 91405 Orsay, France 6 Laboratoire d’Astronomie Spatiale CNRS, Traversée du siphon, les trois lucs, 13120 Marseille, France 7 Centre d’Analyse des Images de l’Institut National des Sciences de l’Univers, CNRS Observatoire de Paris, 61 Avenue de l’Observatoire, 75014 Paris, France

Recent HST determinations of the expansion’s rate of the Universe (the Hubble constant, H_{0}) assumed that the Cepheid period-luminosity relation at V and I are independent of metallicity (Freedman et al., 1996, Saha et al., 1996, Tanvir et al., 1995). The three groups obtain different values for H_{0}. We note that most of this discrepancy stems from the assumption (by both groups) that the periodluminosity relation is independent of metallicity. We come to this conclusion as a result of our study of the period-luminosity relation of 481 Cepheids with 3 million two-colour measurements in the Large Magellanic Cloud and the Small Magellanic Cloud obtained as a by-product of the EROS microlensing survey. We find that the derived interstellar absorption corrections are particularly sensitive to the metallicity, and when our result is applied to recent estimates based on HST Cepheids observations it makes the low-H_{0} values higher and the high-H_{0} value lower, bringing those discrepant estimates into agreement around H_{0} ≈ 70 kms^{−1}Mpc^{−1}. Accepted by: Astronomy & Astrophysics Letters

For preprints, contact beaulieu@astro.rug.nl, sasselov@cfassp36.harvard.edu
Also available from the URL http://xxx.lanl.gov/ps/astro-ph/9612215

Invited Review

Star Formation Histories of Local Group Galaxies

Ludwig-Biermann Award Lecture

Eva K. Grebel^{1,2}

1 Sternwarte der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Astronomisches Institut der Universität Würzburg, Am Hubland, D–97074 Würzburg, Germany

I review the star formation histories of dwarf galaxies in the Local Group. First I investigate which galaxies may be considered members of the Local Group based on various criteria. Out of a sample of 59 galaxies listed in the literature as possible Local Group members, I find the properties of 31 (36) galaxies consistent with likely (potential) Local Group membership. To study the star formation histories of these galaxies, a multi-parameter problem needs to be solved: Ages, metallicities, population fractions, and spatial variations must be determined, which depend crucially on the knowledge of reddening and distance. I summarize the basic methods for studying resolvable stellar populations and illustrate one of the methods using the Fornax dwarf spheroidal galaxy as an example. I present a comprehensive compilation of the star formation histories of dwarf irregulars, dwarf ellipticals, and dwarf spheroidals in the Local Group, and visualize their evolutionary histories through Hodge’s population boxes. All galaxies appear to have differing fractions of old and intermediate-age populations, and those sufficiently massive and undisturbed to retain and recycle their gas are still forming stars today. Star formation has occurred either in distinct episodes or continuously over long periods of time. Metallicities and enrichment vary widely. Constraints on merger and remnant scenarios are discussed, and a unified picture based on the current knowledge is presented. Primary goals for future observations are: accurate age determinations based on turnoff photometry; detection of subpopulations distinct in age, metallicity, and/or spatial distribution; improved distances; and astrometric studies to derive orbits and constrain past and future interactions. Accepted by: Reviews in Modern Astronomy

For preprints, contact grebel@astro.uni-wuerzburg.de
Also available from the URL http://www.astro.uni-wuerzburg.de/~grebel/SFdwarf new.ps.gz
or from the URL http://www.astro.uiuc.edu/~grebel/SFdwarf US.ps.gz